OpenStudy (anonymous):

Solve for x: the quantity of x plus 20 over 2 = 3x I think its x=4

OpenStudy (anonymous):

\[\frac{x+20}{2}=3x\] start with \[x+20=6x\] and solve that one

OpenStudy (anonymous):

yeah you get 4